# Mass gap in non-perturbative quadratic $\mathcal{R}^{2}$ gravity via Dyson-Schwinger

### v1

@@ IV Quantum solution
IV.1 $\Lambda=0$
We can work out the quantum solution exactly in the same way we did for the quartic scalar field in Appendix B. The Dyson-Schwinger equations for our case were obtained in Frasca:2015wva and are the following
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=0,$
$+3G_{2}(x,z)G_{3}(x,x,y)+3G_{2}(x,y)G_{3}(x,x,z)$
$+3G_{2}(x,x)G_{3}(x,y,z)+3G_{1}(x)G_{4}(x,x,y,z)+G_{5}(x,x,x,y,z)=0,$
$+3G_{2}(x,y)G_{4}(x,x,z,w)+3G_{2}(x,z)G_{4}(x,x,y,w)$
$+3G_{2}(x,w)G_{4}(x,x,y,z)+3G_{2}(x,x)G_{4}(x,y,z,w)$
$+3G_{1}(x)G_{5}(x,x,y,z,w)+G_{6}(x,x,x,y,z,w)=0,$
$\vdots$ ,
where we have set
$\mu_{R}^{2}=\frac{M^{2}}{2}-3G_{2}(x,x),$ (41)
for the renormalized mass. In the previous section, we have used $C_{1}$ for classical solutions for the Green’s function, which are not correlation functions as intended in quantum field theory. In the case of the quantum solutions, instead of using $C_{1}$, we have used $G_{2}$ to avoid any further confusion in the present context of the computation. In order to get the Gaussian solution out of this set of equations, we note that all the higher order correlation functions ($n>2$) are expressed through $G_{1}$ and $G_{2}$ but, upon integration for the products $G_{2}(x,x^{\prime})G_{2}(x^{\prime},x)$, we observe that these terms have not overlapping support and yields 0222This is a characteristic of distributional products like $\theta(x)\theta(-x)$ where $\theta(x)$ is the Heaviside function.. This has as a consequence that all the correlation functions for $n>2$, when evaluated at equal points, yield 0, largely simplifying the analysis of the equations and producing an exact solution. Thus, we can use the classical solution for $\Lambda=0$ and write down
$G_{1}(x)=\frac{1}{\sqrt{3}}\mu_{R}\text{dn}\left(p\cdot x+\theta,-1\right),$ (42)
where
$p^{2}=\frac{1}{6}\mu_{R}^{2}.$ (43)
This can be seen as a Fubini-Lipatov solution Fubini:1976jm ; Lipatov:1976ny that breaks translation invariance. For the 2P-correlation function, one has
$G_{2}(p)=\frac{\sqrt{2}\pi^{3}}{K^{3}(-1)}\sum_{n=1}^{\infty}n^{2}\frac{e^{-n\pi}}{1+e^{-2n\pi}}\frac{1}{p^{2}+m_{n}^{2}},$ (44)
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{\mu_{R}}{\sqrt{6}}.$ (45)
From the 2P-correlation function we see that the translation invariance is fully recovered in the propagator and the theory displays a discrete spectrum of excitations with a definite mass gap. Anyway, we have a zero mass particle that is not propagating as it does not contribute to the propagator. The mass term gets a quantum correction by
$G_{2}(x,x)=\int\frac{d^{4}p}{(2\pi)^{4}}G_{2}(p),$ (46)
that diverges and must be renormalized. Then, eq.(41) must be solved self-consistently.
For the Higgs-like solution that preserves translation invariance, we observe a mass renormalization given by eq.(41) but modified by
$\mu_{R}^{2}=M^{2}-6G_{2}(x,x).$ (47)
In this case, the quantum effects could wash out the right sign of the mass term modifying the form of the potential that entails the phase transition.
IV.2 $\Lambda\neq 0$
For this case, we consider the Higgs-like solution. We observe that the cubic equation
$\frac{2\sqrt{2}}{3}M\Lambda-\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi+\phi^{3}=0,$ (48)
has 3 real solutions for
$\frac{3}{16}M^{2}>L.$ (49)
The solutions can be written as
$2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)^{\frac{3}{2}}}\right]-\frac{2}{3}\pi k\right],$ (50)
[…]
$\phi_{0}=2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right].$ (51)
The Dyson-Schwinger set of equations is
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=-\frac{2\sqrt{2}}{3}M\Lambda,$
@@ IV.2 $\Lambda\neq 0$
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-3G_{2}(x,x),$ (54)
that we evaluate in Appendix C. Thus, we have $\phi_{0}$ renormalized by quantum effects. If we write
$M_{R}^{2}=M^{2}-6G_{2}(x,x),$ (55)
we get
$G_{1}=2\sqrt{\frac{1}{3}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)}\cos\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M_{R}\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right].$ (56)
This will yield for the 2P-correlation function
$\partial^{2}G_{2}(x,y)+M_{\Lambda}^{2}G_{2}(x,y)=\delta^{4}(x-y).$ (57)
We have introduced the mass of the particle as
$M_{\Lambda}^{2}=\mu_{R}^{2}\left(4\cos^{2}\left[\frac{1}{3}\arccos\left[\frac{\sqrt{3}M_{R}\Lambda}{2\left(-\frac{4}{3}\Lambda+\frac{1}{2}M_{R}^{2}\right)^{\frac{3}{2}}}\right]-\frac{4}{3}\pi\right]-1\right).$ (58)
It reduces to the previous result for $\Lambda=0$ in eq.(47). Differently form the classical model, the quantum one can be solved exactly.
@@
The scenario here depicted is that of a broken conformal invariance by false vacuum decay through. Correspondingly, a massive particle appears in the theory making the ${\cal R}^{2}$ even more ineffective for the gravitational behavior of the universe. This is true even if the translation invariance is broken by a Fubini-Lipatov instanton for the vacuum as shown for $\Lambda=0$.
In Figure 2, we have plotted the behaviour of the mass gap $M_{\rm\Lambda}$ in a dimensionless unit $M/M_{p}$ with respect to the dimensionless cosmological constant $\Lambda/M^{2}_{p}$. In this plot the blue region is unphysical, as in the regime $\Lambda\gg M^{2}$ which makes the prefactor $\mu^{2}_{R}$ of equation (58) negative. This is clearly visible from equation (54) owing to having an opposite signature in the cosmological constant $\Lambda$ and in the mass gap parameter $M^{2}$, provided $G_{2}(x,x)$ is negligible. On the other hand, in the orange and yellow colored regions, the mass gap $M_{\rm\Lambda}$ increases with the cosmological constant provided the necessary constraint condition $M^{2}\gg\Lambda$ is clearly maintained. In this region due to having this constraint, the prefactor $\mu^{2}_{R}$ of equation (58) is positive. Just like the previous case, when $M^{2}\gg\Lambda$ from equation (54) one can clearly observe that $\mu^{2}_{R}$ is positive. In the right panel of the plot the color coding is helpful in understanding the unphysical as well as the physical region allowed for the present analysis performed in this paper.
@@ V Discussion and Conclusion
  • We have proposed a simple model out of quadratic gravity simplified to the Einstein-Hilbert term plus a quadratic term of the Ricci scalar. This is the Starobinsky model (1). The model is derived assuming that the physics of interest, for strong fields, with conformal symmetry $g_{\mu\nu}=\phi(x)\eta_{\mu\nu}$. Conformal symmetry is broken by quantum effects through the appearance of a mass scale in the quantum theory.
  • The model is well represented through a single mode scalar field characterized by a potential (16) (see Fig. 1) that can display the phenomenon of false vacuum decay, provided a condition is satisfied, already at classical level between the cosmological constant $\Lambda$ and the Starobinsky mass $M$ given by $3M^{2}/16>\Lambda$.
  • The scalar field can be studied and quantized in the limit of a strong field dynamics. For $\Lambda=0$, this yields a mass gap and a spectrum represented in eq.(45). This is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The spectrum depends on $G_{2}(x,x)$, a constant to be renormalized. The mass gap scale is proportional to the Starobinsky mass $M$ when $\Lambda$ is set to 0, as can be foreseen.
@@
  • The quantized theory provides as a solution for the metric, the 1P-correlation function given in eq.(42), breaking translation invariance being equated to a Fubini-Lipatov instanton and representing the vacuum. This symmetry is recovered at the level of 2P-correlation function (44) that describes the propagating degrees of freedom.
  • When the cosmological constant is different from 0, translation invariance is recovered for $G_{1}$ being a constant and the spectrum (58) (see Fig. 2) is characterized by a single excitation, typical of a Higgs-like behavior. Unphysical regions where there is no propagation exist due to the presence of the cosmological constant.
@@ Acknowledgement
Appendix A: GR computations in generalized harmonic coordinates
In this appendix we show how different choices of gauge with harmonic coordinates can give consistent equations of motion for gravity, assuming a conformally flat solution. This is essential for our argument in the main text.
We have to evaluate the Einstein equations
$g^{\delta\gamma}\partial_{\gamma}\partial_{\delta}g_{\alpha\beta}+\partial_{\beta}g^{\gamma\delta}\partial_{\gamma}g_{\alpha\delta}+\partial_{\alpha}g^{\gamma\delta}\partial_{\gamma}g_{\beta\delta}+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}=0,$ (59)
assuming $g_{\alpha\beta}=\phi\eta_{\alpha\beta}$ the above equation becomes
$\phi^{-1}\Box\phi\eta_{\alpha\beta}-\phi^{-2}\partial_{\beta}\phi\eta^{\gamma\delta}\partial_{\gamma}\phi\eta_{\alpha\delta}-\phi^{-2}\partial_{\alpha}\phi\eta^{\gamma\delta}\partial_{\gamma}\phi\eta_{\beta\delta}+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}$
$=\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}=0.$ (60)
Then,
$\Gamma^{\gamma}_{\delta\beta}=\frac{1}{2}g^{\gamma\epsilon}\left[\partial_{\beta}g_{\delta\epsilon}+\partial_{\delta}g_{\beta\epsilon}-\partial_{\epsilon}g_{\delta\beta}\right],$
$\Gamma^{\delta}_{\gamma\alpha}=\frac{1}{2}g^{\delta\rho}\left[\partial_{\alpha}g_{\gamma\rho}+\partial_{\gamma}g_{\alpha\rho}-\partial_{\rho}g_{\gamma\alpha}\right],$ (61)
so that
$\Gamma^{\gamma}_{\delta\beta}=\frac{1}{2}\phi^{-1}\left(\partial_{\beta}\phi\delta^{\gamma}_{\delta}+\partial_{\delta}\phi\delta_{\beta}^{\gamma}-\partial^{\gamma}\phi\eta_{\delta\beta}\right),$
$\Gamma^{\delta}_{\gamma\alpha}=\frac{1}{2}\phi^{-1}\left(\partial_{\alpha}\phi\delta_{\gamma}^{\delta}+\partial_{\gamma}\phi\delta_{\alpha}^{\delta}-\partial^{\delta}\phi\eta_{\gamma\alpha}\right),$ (62)
yielding
$\left(\partial_{\beta}\phi\delta^{\gamma}_{\delta}+\partial_{\delta}\phi\delta_{\beta}^{\gamma}-\partial^{\gamma}\phi\eta_{\delta\beta}\right)\left(\partial_{\alpha}\phi\delta_{\gamma}^{\delta}+\partial_{\gamma}\phi\delta_{\alpha}^{\delta}-\partial^{\delta}\phi\eta_{\gamma\alpha}\right)$
$=4\partial_{\beta}\phi\partial_{\alpha}\phi+\partial_{\beta}\phi\partial_{\alpha}\phi-\partial_{\beta}\phi\partial_{\alpha}\phi+$
$2\partial_{\beta}\phi\partial_{\alpha}\phi-\partial^{\gamma}\phi\partial_{\gamma}\phi\eta_{\alpha\beta}+$
$-\partial_{\beta}\phi\partial_{\alpha}\phi-\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}+\partial_{\alpha}\phi\partial_{\beta}\phi$
$=6\partial_{\alpha}\phi\partial_{\beta}\phi-2\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}.$ (63)
We will get
$\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+2\Gamma^{\gamma}_{\delta\beta}\Gamma^{\delta}_{\gamma\alpha}$
$=\phi^{-1}\Box\phi\eta_{\alpha\beta}-2\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+\phi^{-2}(3\partial_{\alpha}\phi\partial_{\beta}\phi-\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta})$
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi-\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}=0.$ (64)
Now, we introduce
$\phi=e^{f}.$ (65)
This gives
$\Box f\eta_{\alpha\beta}+\partial_{\alpha}f\partial_{\beta}f=0.$ (66)
This equation is not consistent unless $f=constant$. Let us consider the generalized harmonic coordinates then. We will get
$H^{\alpha}=-\phi^{-2}\partial^{\alpha}\phi$ (67)
and then
$H_{\alpha}=-\phi^{-1}\partial_{\alpha}\phi.$ (68)
This gives
$\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha}=2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi$ (69)
and
$H_{\delta}\Gamma^{\delta}_{\alpha\beta}=-\phi^{-1}\partial_{\delta}\phi(\frac{1}{2}\phi^{-1}\left(\partial_{\beta}\phi\delta_{\alpha}^{\delta}+\partial_{\alpha}\phi\delta_{\beta}^{\delta}-\partial^{\delta}\phi\eta_{\alpha\beta}\right))=-\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+\frac{1}{2}\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}.$ (70)
We get
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi+2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi+2\phi^{-2}\partial_{\alpha}\phi\partial_{\beta}\phi-2\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}=0.$ (71)
This equation is inconsistent yet. The only way around is to consider the function $H_{\alpha}$ arbitrary as in in lattice computations Pretorius:2006tp to remove the offending terms. We will have
$\phi^{-1}\Box\phi\eta_{\alpha\beta}+\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi-\phi^{-2}\partial_{\delta}\phi\partial^{\delta}\phi\eta_{\alpha\beta}+(\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha})-2H_{\delta}\Gamma^{\delta}_{\alpha\beta}=0.$ (72)
For consistency reasons we have to chose
$-\phi^{-2}\partial_{\beta}\phi\partial_{\alpha}\phi=(\partial_{\alpha}H_{\beta}+\partial_{\beta}H_{\alpha})+\phi^{-1}\left(\partial_{\beta}\phi H_{\alpha}+\partial_{\alpha}\phi H_{\beta}-H_{\delta}\partial^{\delta}\phi\eta_{\alpha\beta}\right).$ (73)
The choice $\chi/M_{p}=\log\phi$ fixes the gauge properly.
Appendix B: Derivation of Dyson-Schwinger equations for a quartic scalar field
@@
Appendix C: Mass shift
@@
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-3\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+\mu_{R}^{2}},$ (85)
moving to the Euclidean metric in the integral. The integral can be regularized by a cut-off $K$ yielding
$\mu_{R}^{2}=-\frac{4}{3}\Lambda+\frac{M^{2}}{2}-\frac{3}{16\pi^{2}}\left[K^{2}-\mu_{R}^{2}\log\left(1+\frac{K^{2}}{\mu_{R}^{2}}\right)\right].$ (86)
This can be solved by iteration starting with $\mu_{R}=0$ and noticing that, for $M$ very large, the theory has a natural cut-off for $K=M$. Thus, while this correction can appear to be small, it could not be so for the cosmological constant term.

### v2

@@ IV Quantum solution
From the analysis in the precedent section, the true minimum is located at
$G_{1}=\phi_{0}=\frac{1}{6}(3M+\sqrt{3}\sqrt{3M^{2}+8R})\approx\sqrt{\frac{2}{3}R}=\mu_{0},$ (27)
in the considered approximation $3M^{2}<\mathcal{R}$, given the Dyson-Schwinger set of equations
$\partial^{2}G_{1}(x)-\mu_{R}^{2}G_{1}(x)+[G_{1}(x)]^{3}+G_{3}(x,x,x)=\Omega,$
@@
$\mu_{R}^{2}=\mu_{0}^{2}-3G_{2}(x,x),$ (30)
with the correction due to quantum fluctuations evaluated in Appendix B. Thus, we have $\phi_{0}$ renormalized by quantum effects. This will yield for the 2P-correlation function
$\partial^{2}G_{2}(x,y)+M_{G}^{2}G_{2}(x,y)=\delta^{4}(x-y),$ (31)
where the final Mass Gap developed in the theory reads as (see Appendix B)
$M^{2}_{G}=2\mu_{R}^{2}=\frac{8\pi^{2}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}}{27W\left(\frac{16}{27}\pi^{3}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}e^{-1+\gamma+\frac{16\pi^{2}}{3}}\right)},$ (32)
where $W(x)$ is the Lambert function that solves the equation $xe^{x}=w$.
@@
We observe that the conformal factor that appears in passing from the Jordan to the Einstein frame, yields a propagating scalar field. The appearance of a mass gap in the quantum study this Starobinsky scalar field displays a breaking of the scale symmetry, reducing the geometry of the spacetime to that of the backbone geometry that has a null Ricci curvature. This shows a transition between a full blown quadratic gravity to a Einstein-Hilbert dominated gravity.
[…]
@@ V Discussion and Conclusion
  • We have proposed a simple model out of quadratic gravity simplified to the Einstein-Hilbert term plus a quadratic term of the Ricci scalar. This is the Starobinsky model (1). The model is derived assuming that the physics of interest, for strong fields, is described by a large and constant Ricci scalar granting the appearance of a mass gap in the Starobinsky scalaron field. This provides a transition scale from a quadratic dominated gravity to a Einstein-Hilbert dominated gravity.
  • The model is well represented through a single mode scalar field characterized by a potential (11) (see Fig. 1) that can display the phenomenon of false vacuum decay, provided a condition is satisfied, already at classical level between the scalar Ricci curvature, that we assume constant, $\mathcal{R}$ satisfies the strong coupling condition $3M^{2}<\mathcal{R}$, being $M$ the mass term due to the quadratic term in the action (1).
  • The scalar field can be studied and quantized in the limit of a strong field dynamics. For $\Lambda=0$, this yields a mass gap and a spectrum represented in eq.(22). This is a discrete spectrum with an infinite tower of harmonic-oscillator like excitations. The spectrum depends on $G_{2}(x,x)$, a constant to be renormalized. The mass gap scale is proportional to the Ricci scalar $\mathcal{R}$.
@@
  • The quantized theory provides as a solution for the metric, the 1P-correlation function given in eq.(27), and representing the vacuum of the theory. A mass gap is displayed at the level of 2P-correlation function (31) that describes the propagating degrees of freedom.
@@
Similarly, an interesting area of application of our formalism is for Einstein-Gauss-Bonnet cosmological models Nojiri:2005jg ; Cruz-DombrizS:2012 ; BenettiSCAL:2018 ; OdintsovOFF:2020 ; DeLaurentis:2015fea that appear very promising in the understanding of the experimental data Odintsov:2025kyw . We will address also this matter in future publications.
@@ Acknowledgement
Appendix A: Derivation of Dyson-Schwinger equations for a quartic scalar field
@@
Appendix B: Mass shift
@@
$\mu_{R}^{2}=\mu_{0}^{2}-3\int\frac{d^{4}p}{(2\pi)^{4}}\frac{1}{p^{2}+2\mu_{R}^{2}}.$ (44)
Using dimensional regularization, we get
$\int\frac{d^{D}p}{(2\pi)^{4}}\frac{1}{p^{2}+2\mu_{R}^{2}}=\frac{\Gamma(-1+\epsilon)}{(4\pi)^{2-\epsilon}}(\mu_{R}^{2})^{1-\epsilon},$ (45)
where we set $\epsilon=(4-D)/2$. Thus, doing an expansion for $\epsilon\rightarrow 0$, we get
$\mu_{R}^{2}=\mu_{0}^{2}-3\left[-\frac{\mu_{R}^{2}}{16\pi^{2}\epsilon}+\frac{\mu_{R}^{2}}{16\pi^{2}}\left(-1+\gamma+\log(\mu_{R}^{2})+\log(4\pi)\right)\right]+O(\epsilon),$ (46)
where $\gamma$ is the Euler-Mascheroni constant. We take care of the divergent part by considering it a correction to the coupling constant $M$. Thus, we can safely take the limit $\epsilon\rightarrow 0$ and we are left with
$\mu_{R}^{2}\left[1+\frac{3}{16\pi^{2}}\left(-1+\gamma+\log(\mu_{R}^{2})+\log(4\pi)\right)\right]=\mu_{0}^{2}.$ (47)
This equation can be solved using the Lambert function $W$, writing
$\mu_{R}^{2}=\frac{4\pi^{2}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}}{27W\left(\frac{16}{27}\pi^{3}\left(\sqrt{3}\sqrt{3M^{2}+8R}+3M\right)^{2}e^{-1+\gamma+\frac{16\pi^{2}}{3}}\right)}.$ (48)